# Why machine learning models fail to fully capture epistemic uncertainty

### v1

@@ 4.3 Synthetic data illustration
To illustrate the limitations of existing uncertainty quantification methods in producing reliable disentangled estimates, we showcase the behavior of the eight methods described in Section 4.1 for the problem setting described in Example 3.2. This section focuses primarily on the results obtained using Deep Ensembles, but we emphasize that the results for the other seven methods are also consistent with Deep Ensembles, and coherent with our argumentation. Results for the remaining seven methods are provided in Appendix B, and detailed implementation specifications are included in Appendix A.2.
@@
Figure 4 shows a detailed illustration of the results for Deep Ensembles. Although the aleatoric uncertainty estimates on the right-hand side of the data support improve as the sample size increases, the behaviour on the left-hand side remains problematic. Specifically, the model persistently overestimates aleatoric uncertainty. A plausible explanation is that the approximate symmetry of the observations around the learned conditional mean induces the model to misinterpret systematic deviation as inherent data noise. Consequently, this misestimation is not corrected with additional data; instead, the model becomes increasingly confident in its erroneous predictions and continues to interpret the discrepancy from the ground-truth conditional mean as aleatoric variability. By contrast, epistemic uncertainty decreases monotonically with sample size, mirroring the behaviour observed for the reference distribution. The often-repeated claim in the literature that aleatoric uncertainty estimates can be trusted once epistemic uncertainty is low is therefore not correct: when the training set is sufficiently large, in a difficult-to-learn setting, the model remains overconfident in highly biased predictions, producing inflated aleatoric uncertainty estimates. These observations support our two principal conclusions. First, current measures of epistemic uncertainty and existing disentanglement frameworks fail to capture systematic bias, thereby underestimating epistemic uncertainty. Second, the resulting aleatoric-uncertainty estimates misclassify epistemic uncertainty with data noise and hence do not achieve a precise disentanglement of the two sources of uncertainty. Bias is approximated as specified in Eqn. 2.2: datasets are repeatedly resampled from the true data-generating process, models are trained on each sample, and the mean deviation of the predictions from the true conditional mean is computed. Comprehensive results for every method are presented in Table 1; the table shows that bias is frequently larger than—and largely uncorrelated with—the corresponding epistemic uncertainty estimates. These findings underscore our central position: Any attempt to disentangle aleatoric from epistemic uncertainty must assess not only how each component is modelled, but also the extent to which systematic bias distorts their partitioning.
The results for the remaining methods are provided in Appendix A. In this section, we highlight two additional key observations. First, we observe that Bayesian approaches—specifically Laplace Approximation (LA) and Hamiltonian Monte Carlo (HMC)—require a larger amount of data to produce accurate estimates of aleatoric uncertainty in the right-hand region of the input space. This behavior aligns with theoretical expectations of Bayesian inference: the increased data requirement reflects the greater difficulty these models encounter in learning the underlying data-generating function, which in turn leads to higher epistemic uncertainty—consistent with models that make inaccurate predictions. Second, the Deep Evidential Regression (DER) method consistently produces poor estimates of aleatoric uncertainty across all sample sizes and throughout the entire input domain. These inaccuracies are accompanied by overconfident predictions, as evidenced by the systematic underestimation of epistemic uncertainty. This flawed characterization of uncertainty is consistent with findings reported in recent literature [17].
@@ 5 Discussion
In this work, we address the problem of blurriness in the definition of epistemic uncertainty for machine learning models and the difficulties that come with it. In attempting to solve this problem, we first unified different definitions of epistemic uncertainties for machine learning models into a formal taxonomy. In our illustrative simulations, we work within a heteroscedastic regression framework; nevertheless, the key findings are model-independent and hence translate to classification problems. Using a statistical sampling-based evaluation protocol, we showed that even if parts of the arising epistemic uncertainty are well captured, the underlying model bias can still lead to an overall underestimate of epistemic uncertainty. By differentiating between procedural and data uncertainty, we find that different uncertainty estimation methods merely capture parts of the remaining epistemic uncertainty. Furthermore, the non-assessment of bias seemingly leads to an artificially high estimate of aleatoric uncertainty. Our call is therefore not only for deriving more task-specific and fine-grained evaluation protocols of epistemic uncertainty in practical applications, but also a proper analysis of model bias when estimating it. It might be argued that the goal of an epistemic uncertainty estimate is not to quantify bias. However, precisely by not doing so, the disentanglement of uncertainties becomes problematic.
@@ Appendix A Experimental details
A.1 Reference distribution approximation
To obtain an empirical approximation of the reference distribution, we sample $20$ different datasets. For each dataset, we train $10$ different MLP models with different random weight initialization. Each MLP uses the whole dataset. The MLP models have $4$ hidden layers with $100$ neurons each. We use a learning rate of 0.009 and train the models for $500$ epochs. We repeated the process for different training dataset sizes: $50,100,500$ with batch sizes $32,32,64$ respectively. Seeds for the datasets are $\{7,42,2024,123,999,50,100,150,200,250,300,350,400,450,500,550,600,650,700,750\}$.
A.2 Experimental details by method
All methods are evaluated for different training dataset sizes: $\{50,100,500\}$ with batch sizes $\{32,32,64\}$ respectively. The experiments are repeated for $4$ different seeds: $7,42,123,999$ and $2024$. To conduct the experiments, a hyperparameter tuning over relevant parameters was performed.
@@ A.2 Experimental details by method
Appendix B Further experimental results
In this section, we include further experimental results for the other methods that we tested. The data is generated according to Example 3.2. We conduct the same experiment as described for the Deep Ensemble case in Section 4.3. The hyperparameters of each method were chosen according to Section A. The results can be seen in Figure 5-11. We see that every method with a low epistemic uncertainty estimate overestimates the aleatoric uncertainty in the left-hand side of the figure due to their inability to acknowledge the bias as part of their epistemic uncertainty representations. Also, for more data-intensive methods, such as the Bayesian approaches, a more careful and realistic estimate of the epistemic uncertainty is obtained.
@@ Appendix B Further experimental results
Training sample size: N = 100 Model Aleatoric unc. Epistemic unc. Bias $\hat{\sigma}(x)-\sigma(x)$ Reference 0.83 (0.01) 0.041 (0.005) 0.084 (0.002) 0.248 (0.004) Bootstrap 0.87 (0.167) 0.006 (0.008) 0.125 (0.022) 0.239 (0.027) DeepEnsemble 0.87 (0.148) 0.007 (0.011) 0.147 (0.018) 0.235 (0.021) EDL 0.21 (0.002) 0.003 (0.0) 0.269 (0.01) 0.751 (0.005) ViELBONet 0.95 (0.064) 0.079 (0.001) 0.127 (0.021) 0.234 (0.022) MCDropoutNet 0.9 (0.097) 0.009 (0.002) 0.19 (0.026) 0.334 (0.033) HMCNet 9.08 (13.734) 2.131 (0.224) 0.179 (0.015) 7.87 (0.853) LaplaceNet 2.43 (5.559) 4.211 (3.047) 0.133 (0.024) 0.8 (0.54) GP 0.86 (0.048) 0.008 (0.001) 0.194 (0.013) 0.308 (0.02)
Training sample size: N = 500 Model Aleatoric unc. Epistemic unc. Bias $\hat{\sigma}(x)-\sigma(x)$ Reference 1.04 (0.016) 0.013 (0.001) 0.044 (0.001) 0.076 (0.003) Bootstrap 0.98 (0.1) 0.003 (0.004) 0.085 (0.001) 0.131 (0.01) DeepEnsemble 0.99 (0.136) 0.004 (0.005) 0.1 (0.002) 0.122 (0.01) EDL 0.67 (0.247) 0.013 (0.007) 0.168 (0.016) 0.967 (0.031) ViELBONet 1.06 (0.115) 0.009 (0.0) 0.097 (0.005) 0.122 (0.012) MCDropoutNet 1.06 (0.079) 0.008 (0.003) 0.119 (0.006) 0.163 (0.008) HMCNet 4.17 (20.356) 0.378 (0.017) 0.129 (0.003) 1.936 (1.495) LaplaceNet 1.29 (3.777) 0.487 (3.802) 0.115 (0.016) 0.187 (0.345) GP 1.04 (0.069) 0.008 (0.003) 0.172 (0.005) 0.152 (0.014)

### v2

@@ 4.4 Synthetic data illustration
To illustrate the limitations of existing uncertainty quantification methods in producing reliable disentangled estimates, we showcase the behavior of the eight methods described in Section 4.1 for the problem setting described in Section 2. This section focuses primarily on the results obtained using Deep Ensembles, but we emphasize that the results for the other seven methods are also consistent with Deep Ensembles and coherent with our argumentation. Results for the remaining seven methods and detailed implementation specifications are included in Appendix A.1.
@@
Figure 3 illustrates that epistemic uncertainty estimates decrease monotonically with sample size, consistent with the behavior observed for the reference distribution. This trend confirms that, as the dataset grows, models become increasingly confident, yet not necessarily more accurate, particularly in difficult-to-learn regions. In these settings, even with abundant data, the model can remain overconfident in systematically biased predictions. Such behavior reveals a key limitation of current epistemic uncertainty measures and disentanglement frameworks: they fail to account for systematic bias, leading to an underestimation of epistemic uncertainty. This pattern is not unique to Deep Ensembles, but is consistently observed across other methods (see Figures 814), indicating that this limitation is intrinsic to the way epistemic uncertainty is commonly represented rather than a method-specific artifact. Bias is approximated as specified in Eqn. 2.2: datasets are repeatedly resampled from the true data‐generating process, models are trained on each sample, and the mean deviation of the predictions from the true conditional mean is computed. Comprehensive results for every method are presented in Table 1; the table shows that bias is frequently larger than –and largely uncorrelated with– the corresponding epistemic uncertainty estimates. These findings underscore our central position: any attempt to disentangle aleatoric from epistemic uncertainty must assess not only how each component is modeled, but also the extent to which systematic bias distorts their partitioning.
Figures 6(c) and 7 further illustrate the relationship between the epistemic uncertainty estimates and the procedural and data uncertainty components obtained from the reference distribution. When comparing these components, the relation shifts upward once data uncertainty is included alongside procedural uncertainty, indicating that most methods primarily capture the procedural component of epistemic uncertainty. This observation highlights the necessity of adopting a more fine-grained taxonomy to identify which parts of epistemic uncertainty are being represented, as well as the importance of the reference distribution in empirically validating these estimates. On the other hand, the aleatoric uncertainty estimates on the right‐hand side of the data support improve as the sample size increases, the behaviour on the left‐hand side remains problematic. Specifically, the model persistently overestimates aleatoric uncertainty. A plausible explanation is that the approximate symmetry of the observations around the learned conditional mean induces the model to misinterpret systematic deviation as inherent data noise. Consequently, this misestimation is not corrected with additional data; instead, the model becomes increasingly confident in its erroneous predictions and continues to interpret the discrepancy from the ground-truth conditional mean as aleatoric uncertainty. This behavior is illustrated in Figure 5(a), which compares the average estimated aleatoric uncertainty with the true aleatoric noise in regions of low and high model bias. The figure shows that in areas of high bias, the estimated aleatoric uncertainty diverges substantially from the true value, demonstrating that biased predictions systematically distort the disentanglement of aleatoric and epistemic uncertainty.
4.5 Real data experiments
In this section, we extend our analysis to the New York taxi trip duration dataset published in this Kaggle competition. We begin by describing how the reference distribution is estimated, followed by an explanation of how we approximate the ground-truth aleatoric uncertainty. Finally, we present the results obtained with the Deep Ensemble method, and refer to the appendix for the results of the rest of the methods. The dataset consists of taxi trip records from New York City, where the target variable corresponds to the trip duration and the feature space comprises 14 input variables. The training set contains $1,206,737$ observations, and the test set includes $230,406$ observations. Details about the dataset source and preprocessing steps are provided in Appendix A.2. This dataset is particularly suitable for evaluating uncertainty estimation methods for several reasons. First, its large sample size allows us to analyze how epistemic uncertainty behaves when the data-driven component becomes relatively small compared to the model and procedural components, allowing us to estimate data and procedural uncertainty via the reference distribution. In the same sense, a simulated estimate of ground truth aleatoric uncertainty can be obtained as we explain later. Second, since the task is formulated as a heteroscedastic regression problem, where trip durations exhibit high variability depending on factors such as traffic conditions or route choice, it provides a natural setting for disentangling aleatoric from epistemic uncertainty. Finally, the dataset’s scale and heterogeneity make it a realistic benchmark for assessing whether uncertainty quantification methods can remain well-calibrated and informative in complex, high-dimensional environments. As in the synthetic data experiments, we assume that the target variable follows a normal distribution, and the model outputs both the mean and variance of this distribution.
Reference distribution. To estimate the reference distribution as described in Section A.1, we partition the training dataset into $100$ disjoint subsets of $12,067$ observations each (i.e., $n_{d}=100$). For each subset, we train $n_{\gamma}=5$ multilayer perceptrons (MLPs) with different random weight initializations, resulting in a total of $500$ independently trained models. Each model produces predictive means $\hat{y}_{N,\gamma}(\boldsymbol{x})$, and by computing the empirical variance across these 500 predictions, we obtain an estimate of $\text{Var}_{\mathcal{D}_{N},\gamma}(\hat{y}|\boldsymbol{x})$ as defined in Eqn. 7. This procedure operationalizes the reference distribution introduced in Section 3: the pseudo-resampling across datasets ($\mathcal{D}_{N}$) captures data uncertainty while the random initializations account for procedural uncertainty. Given the large sample size of the dataset, the partitioning approach provides a computationally feasible approximation of repeated sampling from the underlying data-generating process, effectively emulating access to its stochastic variability. This design allows us to decompose the epistemic uncertainty component into its data and procedural parts in a controlled manner, while being in a real world data setup. Figure 4(c) shows the resulting decomposition of total epistemic variance into procedural and data components. We observe that procedural uncertainty dominates across most of the input domain. This outcome reflects the high capacity and non-convex optimization landscape of neural networks: even with abundant data, different random initializations can converge to distinct local minima, leading to substantial procedural variation. By contrast, data uncertainty is smaller, indicating that resampling subsets from such a large dataset induces only limited diversity in the fitted models. It is important to note that the relative magnitudes of these two components depend on the experimental configuration. Increasing the subset size $n_{d}$ would typically reduce data uncertainty, while enlarging $n_{\gamma}$ would yield a more precise estimate of procedural uncertainty.
Aleatoric uncertainty. To approximate the ground-truth aleatoric uncertainty, we train an ensemble of $n_{\gamma}=5$ multilayer perceptrons (MLPs) on the full training dataset of $1,206,737$ observations. Since aleatoric uncertainty represents the irreducible noise inherent to the data-generating process, training on the complete dataset minimizes the influence of data-driven and procedural components, yielding the most reliable estimate of this term attainable in practice. We take the mean predicted variance across the ensemble members as the reference estimate of aleatoric uncertainty. To assess the effect of limited data and model bias on uncertainty disentanglement, we then compare this reference estimate with those obtained when each model is trained using only $1\%$ of the available data. This comparison highlights how insufficient data coverage and higher epistemic uncertainty can distort the estimated aleatoric component, revealing the extent to which current methods conflate epistemic and aleatoric sources of uncertainty.
Deep Ensemble estimates. As discussed in Example 1 and in the synthetic data experiments, Deep Ensembles predominantly capture the procedural component of epistemic uncertainty. This limitation arises because ensemble diversity originates solely from random initialization and stochastic training dynamics, while all models are trained on the same dataset. Consequently, the variance across ensemble members reflects the sensitivity of the optimization process rather than uncertainty about the data distribution itself. This behavior is observed in Figure 4(b), where the total variance of the reference distribution, represented on the $y$-axis as the sum of data and procedural uncertainty, is systematically higher than the epistemic uncertainty predicted by the Deep Ensemble. Quantitatively, the epistemic uncertainty estimate underestimates the reference distribution variance in approximately $80\%$ of the predictive samples. This systematic underestimation indicates that the ensemble fails to account for the data-driven component of epistemic uncertainty, which captures how model predictions would vary under different training datasets drawn from the same data-generating process. To further examine this relationship, Figure 4(a) decomposes the reference distribution into its procedural and data uncertainty components. When comparing only the procedural component of the reference distribution with the epistemic uncertainty predicted by the Deep Ensemble, the discrepancy disappears. This alignment confirms that Deep Ensembles primarily quantify procedural uncertainty, that is, the uncertainty arising from stochasticity in optimization and model initialization, while largely neglecting the variability induced by finite data sampling.
In practical terms, this result reinforces the theoretical argument presented in Section 4.3: the epistemic uncertainty estimated by Deep Ensembles does not reflect the full reducible uncertainty in the model but only its procedural subcomponent. Consequently, downstream tasks that rely on a correct epistemic uncertainty estimation, such as active learning, model selection, or out-of-distribution detection, may be misled by an incomplete uncertainty representation. In the same sense, Figure 5(b) shows that in the region where the model is more incorrect, the difference between the estimated aleatoric uncertainty and the simulated ground truth aleatoric uncertainty is larger than in the region where the model is more accurate, showing how a high bias can distort the disentanglement of aleatoric and epistemic uncertainty. The findings from this real-world regression task, therefore, underscore the importance of distinguishing between procedural and data uncertainty when evaluating or designing second-order uncertainty quantification methods.
Figure 4: (a): Procedural and data uncertainty estimates from the reference distribution of the taxi trip duration dataset. The procedural uncertainty component is larger than the data uncertainty component. (b): Comparison between the epistemic uncertainty estimated by the Deep Ensemble and the epistemic uncertainty obtained from the reference distribution. The Deep Ensemble underestimates the epistemic uncertainty $80\%$ of the time. (c): Relation between the epistemic uncertainty estimated by the Deep Ensemble and the procedural uncertainty coming from the reference distribution. On average the Deep Ensemble’s estimate captures the procedural uncertainty part of the total epistemic uncertainty.
@@ 5 Discussion
Epistemic uncertainty is generally not well-defined and still very much a heuristic concept for machine learning models, leading to interpretability issues. As a first attempt to formalise a more fine-grained notion, we present a taxonomy that unifies the leading definitions of epistemic uncertainty in the literature, and present a simulation-based evaluation protocol for it. Although our illustrative simulations involve heteroscedastic regression problems, our key findings are also applicable to homoscedastic regression and classification settings. Our simulations include one synthetic data experiment and one real data experiment. Whilst the synthetic data experiment allows for resampling, bias calculation, and comparison with the true aleatoric uncertainty, the taxi data experiment requires some modeling design decisions, which might introduce bias in the result. With these experiments, we show that epistemic uncertainty estimates underestimate the uncertainty by capturing only some part of one of the components described in our taxonomy, and ignoring the model bias. More precisely, the evaluated methods seem to capture the procedural component, while being away from the data uncertainty. This underestimation also induces an overestimation of aleatoric uncertainty in regions where the predictive bias is larger. An accurate estimation of the epistemic uncertainty requires deriving more task-specific and fine-grained evaluation protocols of the estimates in practical applications, and assessing the model bias when estimating it. It might be argued that the goal of an epistemic uncertainty estimate is not to quantify bias. However, precisely by not doing so, the disentanglement of uncertainties becomes problematic. Our experiments show how the aleatoric uncertainty estimates can become largely incorrect in regions where the model error increases, undermining the disentanglement of the two uncertainties.
@@ Appendix A Experimental details
A.1 Synthetic data experiments
Reference distribution approximation
To obtain an empirical approximation of the reference distribution, we sample $20$ different datasets. For each dataset, we train $10$ different MLP models with different random weight initializations. Each MLP uses the whole dataset. The MLP models have $4$ hidden layers with $100$ neurons each. We use a learning rate of 0.009 and train the models for $500$ epochs. We repeated the process for different training dataset sizes: $50,100,500$ with batch sizes $32,32,64$ respectively. Seeds for the datasets are $\{7,42,2024,123,999,50,100,150,200,250,300,350,400,450,500,550,600,650,700,750\}$.
Experimental details by method
All methods are evaluated for different training dataset sizes: $\{50,100,500\}$ with batch sizes $\{32,32,64\}$ respectively. The experiments are repeated for $5$ different seeds: $7,42,123,999$ and $2024$. To conduct the experiments, a hyperparameter tuning over relevant parameters was performed.
@@ A.1 Synthetic data experiments
Further experimental results
In this section, we present additional experimental results for the other methods we tested. The data is generated according to the setting described in Section 2. We conduct the same experiment as described for the Deep Ensemble case in Section 4.4. The hyperparameters of each method were chosen according to Section A. The results can be seen in Figure 8-14. We highlight two additional key observations: First, we observe that Bayesian approaches –specifically Laplace Approximation (LA) and Hamiltonian Monte Carlo (HMC)– require a larger amount of data to produce accurate estimates of aleatoric uncertainty in the right-hand region of the input space. This behavior aligns with theoretical expectations of Bayesian inference: The increased data requirement reflects the greater difficulty these models encounter in learning the underlying data-generating function, which in turn leads to higher epistemic uncertainty–consistent with models that make inaccurate predictions. Second, the Deep Evidential Regression (DER) method consistently produces poor estimates of aleatoric uncertainty across all sample sizes and throughout the entire input domain. These inaccuracies are accompanied by overconfident predictions, as evidenced by the systematic underestimation of epistemic uncertainty. This flawed characterization of uncertainty is consistent with findings reported in recent literature [24]. We see that every method with a low epistemic uncertainty estimate overestimates the aleatoric uncertainty in the left-hand side of the figure due to their inability to acknowledge the bias as part of their epistemic uncertainty representations. Also, for more data-intensive methods, such as the Bayesian approaches, a more careful and realistic estimate of the epistemic uncertainty is obtained.
(a) Synthetic data (Equation 8)
(b) Taxi dataset.
Figure 5: Comparison of the estimated aleatoric uncertainty with the true (simulated) aleatoric noise across regions characterized by different levels of model bias and inherent data noise. The notation $\sigma_{i}$ for $i\in{1,2,3,4}$ denotes the true aleatoric uncertainty for each of the four regions: 1. low bias and low true noise, 2. high bias and low true noise, 3. low bias and high true noise, and 4. high bias and high true noise. For the synthetic dataset, the division between high- and low-noise regions follows Eqn. 8: points with $x<0.6$ are considered low-noise, while those with $x\geq 0.6$ are considered high-noise. Within each noise region, samples are ranked according to their model bias, and the lower half of points (by bias magnitude) defines the low-bias region, while the upper half defines the high-bias region. For the taxi dataset, the high- and low-noise regions are determined using the $90$-th percentile of the simulated true aleatoric uncertainty, as described in Section 4.5. Within each noise region, data points are further divided into low- and high-bias subsets based on the 70th percentile of the estimated model bias. This construction allows for a direct comparison between the estimated and true aleatoric uncertainty across systematically defined regions, highlighting how model bias affects the reliability of aleatoric uncertainty estimates in both synthetic and real-world scenarios.
Figure 6: Epistemic uncertainty estimates for the synthetic dataset. (a): Decomposition of the epistemic uncertainty estimated with the reference distribution into procedural and data uncertainties. There exist a strong correlation between both, which is expected, and procedural uncertainty overweighs across almost all the data points. (b): Comparison between the epistemic uncertainty estimated with the reference distribution (both data and procedural components) and the epistemic uncertainty estimated with the Deep Ensemble. Here, the epistemic part from the reference distribution is larger, showing that the Deep Ensemble underestimates this uncertainty. (c): When the data uncertainty component is removed from the estimate of the reference distribution, both estimates are much closer, suggesting that the Deep Ensemble fails to capture the data uncertainty component of the epistemic uncertainty.
(a) Bootstrap
(b) Bootstrap
(c) GP
(d) GP
(e) EDL
(f) EDL
(g) Laplace
(h) Laplace
(i) MC Dropout
(j) MC Dropout
(k) VI
(l) VI
Figure 7: Comparison of procedural and data uncertainties vs epistemic uncertainty across methods on the Sine dataset.
@@
Training sample size: $N=100$
Model $\hat{\sigma}(x)$ $\text{Var}\big[\hat{y}|\boldsymbol{x}\big]$ $\mathbb{E}_{\mathcal{D}_{N}}[\hat{y}(x)-f(x_{i})]$ $\hat{\sigma}(x)-\sigma(x)$
Reference 0.83 (0.01) 0.041 (0.005) 0.084 (0.002) 0.248 (0.004)
Bootstrap 0.87 (0.167) 0.006 (0.008) 0.125 (0.022) 0.239 (0.027)
DeepEnsemble 0.87 (0.148) 0.007 (0.011) 0.147 (0.018) 0.235 (0.021)
EDL 0.21 (0.002) 0.003 (0.0) 0.269 (0.01) 0.751 (0.005)
ViELBONet 0.95 (0.064) 0.079 (0.001) 0.127 (0.021) 0.234 (0.022)
MCDropoutNet 0.90 (0.097) 0.009 (0.002) 0.190 (0.026) 0.334 (0.033)
HMCNet 9.08 (13.734) 2.131 (0.224) 0.179 (0.015) 7.870 (0.853)
LaplaceNet 2.43 (5.559) 4.211 (3.047) 0.133 (0.024) 0.800 (0.540)
GP 0.86 (0.048) 0.008 (0.001) 0.194 (0.013) 0.308 (0.020)
Training sample size: $N=500$
Model $\hat{\sigma}(x)$ $\text{Var}\big[\hat{y}|\boldsymbol{x}\big]$ $\mathbb{E}_{\mathcal{D}_{N}}[\hat{y}(x)-f(x_{i})]$ $\hat{\sigma}(x)-\sigma(x)$
Reference 1.04 (0.016) 0.013 (0.001) 0.044 (0.001) 0.076 (0.003)
Bootstrap 0.98 (0.100) 0.003 (0.004) 0.085 (0.001) 0.131 (0.010)
DeepEnsemble 0.99 (0.136) 0.004 (0.005) 0.100 (0.002) 0.122 (0.010)
EDL 0.67 (0.247) 0.013 (0.007) 0.168 (0.016) 0.967 (0.031)
ViELBONet 1.06 (0.115) 0.009 (0.000) 0.097 (0.005) 0.122 (0.012)
MCDropoutNet 1.06 (0.079) 0.008 (0.003) 0.119 (0.006) 0.163 (0.008)
HMCNet 4.17 (20.356) 0.378 (0.017) 0.129 (0.003) 1.936 (1.495)
LaplaceNet 1.29 (3.777) 0.487 (3.802) 0.115 (0.016) 0.187 (0.345)
GP 1.04 (0.069) 0.008 (0.003) 0.172 (0.005) 0.152 (0.014)
@@
A.2 Real data experiments
Data preprocessing
The raw dataset used for this task can be found in this Kaggle competition. The goal is to predict the duration of a taxi trip in New York. The dataset contains $1,458,644$ observations from 2016. The code to preprocess the data is in the GitHub repository of this paper. The target variable is transformed according to $y^{\prime}=\log(y+1)$.
Data splitting. We consider the data before June as the training dataset and the data after June as the test dataset. In this way, we get $1,206,737$ training points and $230,406$ test points. However, only $1\%$ of the dataset is used for training the models. The simulated ground truth aleatoric uncertainty is obtained by using the complete dataset as explained in Section 4.
Models
Deep Ensemble. Trained with an MLP of $10$ hidden layers with $300$ neurons each, learning rate of $0.0009$. Training epochs: $50$. Every member of the deep ensemble uses the whole dataset (containing $1\%$ of the total training dataset).
Bootstrap Ensemble. Trained as the Deep Ensemble, but each ensemble member uses a boostrap sample with $60\%$ of the training data (out of the $1\%$).
Monte Carlo Dropout. A single MLP of $10$ hidden layers with $300$ neurons each is trained. A dropout rate of $0.2$ is used at training and inference time. A learning rate of $0.0009$ is used.
Variational Inference. This model is trained based on Blundell et al. [52]. An MLP of $10$ hidden layers with $300$ neurons each converted to a Bayesian neural network is used. The model is trained with a learning rate of $0.0009$, 2020 burn-in epochs, a beta value of $10$, $10$ Monte Carlo samples at training time, and $500$ at test. The model is evaluated for $50$ training epochs.
Laplace Approximation. An MLP of $4$ hidden layers with $100$ neurons each is used to find the MAP estimate, together with a learning rate of $0.0009$. For the Laplace step, a prior precision of $1$ and a noise value of $0.1$ are used. The model is trained for $50$ epochs, and uses $100$ samples from the posterior distribution.
Hamiltonian Monte Carlo. An MLP of $10$ hidden layers with $300$ neurons each is used as the underlying model. The sampling step consists of $70$ samples with a sampling step of $0.00015$. The first $10$ samples are always burnt. A $\tau$ value of $1$ is used. At inference time, we sample $100$ times from the posterior distribution and drop the first $20$. The model is trained for $50$ epochs.
Heteroscedastic Gaussian Process. The implementation is the same as in the synthetic data case. The number of inducing points (see [20] for an explanation) is set to $256$, the learning rate to $0.0009$, epochs to $50$.
A.2.1 Further experimental results
In Section 4.5 we discuss how the Deep Ensembles capture only the procedural part of the epistemic uncertainty, effectively underestimating it. In Figure 15 we can see how the Bootstrap and Monte Carlo Dropout methods have a very similar behavior: a strong correlation between the epistemic uncertainty and the procedural uncertainty, that is shifted when the data uncertainty component is considered.
Model $\hat{\sigma}^{2}$ $\hat{\sigma}^{2}(x)-\sigma_{*}^{2}(x)$ $\mathrm{Var}\!\big[\hat{y}|\boldsymbol{x}\big]$ $\mathrm{Var}_{\mathcal{D}_{N},\gamma}\!\big(\hat{y}|\boldsymbol{x}\big)$- $\mathrm{Var}\!\big[\hat{y}|\boldsymbol{x}\big]$ $\mathbb{E}_{\mathcal{D}_{N}}\!\big[\mathrm{Var}_{\gamma}\!\big(\hat{y}|\boldsymbol{x}\big)|\mathcal{D}_{N}\big]$- $\mathrm{Var}\!\big[\hat{y}|\boldsymbol{x}\big]$
DE 0.1222 0.0004 0.0083 0.0046 0.0013
Bootstrap 0.1127 0.0091 0.0082 0.0047 0.0014
MC Dropout 0.1284 0.0066 0.0106 0.0023 0.001
VI 1.8859 1.7641 0.0055 0.0074 0.0041
Laplace 0.1754 0.0536 0.0137 0.0008 0.0041
HMC 0.2319 0.1101 0.1148 0.1019 0.1052
GP 1.0334 0.9116 0.0636 0.0507 0.054
Table 2: Taxi dataset uncertainties. $\text{Var}\big[\hat{y}|\boldsymbol{x}\big]$ represents the epistemic uncertainty estimated by each model. $\text{Var}_{\mathcal{D}_{N},\gamma}(\hat{y}|\boldsymbol{x})$ corresponds to the epistemic uncertainty obtained with the reference distribution, while $\mathbb{E}_{\mathcal{D}_{N}}\big[\text{Var}_{\gamma}(\hat{y}|\boldsymbol{x})|\mathcal{D}_{N}\big]$ is the procedural uncertainty obtained by the reference distribution. The values are the average over all test data points.
(a) Bootstrap
(b) Bootstrap
(c) GP
(d) GP
(e) HMC
(f) HMC
(g) Laplace
(h) Laplace
(i) MC Dropout
(j) MC Dropout
(k) VI
(l) VI
Figure 15: Comparison of procedural and data uncertainties vs epistemic uncertainty across methods on the Taxi dataset. Figures in purple (first figure of every method) show the relation between the epistemic uncertainty from the reference distribution (including procedural and data uncertainties), and green pictures (second figure of every method) show the relationship between the procedural uncertainty from the reference distribution and the epistemic uncertainty from each method. In general, we see that adding the data uncertainty component shifts the relation between the two estimates.